# Statistically-Guided Dual-Domain Meta-Learning with Adaptive Multi-Prototype Aggregation for Distributed Fiber Optic Sensing

### v1

@@ 5.3 Comparison Experiments by Category
Per-class accuracy highlights the performance of each method in cross-deployment testing.
On the OSDG1 dataset, the results in Table 6 indicate that many baseline methods suffer from severe class skew. The 2D-CNN achieved near-perfect accuracy on the Background, Climb, and Rain classes, but its performance on the Smash class was extremely poor. Conversely, KNN demonstrated high accuracy on the Rain class, achieving 0.9919, but performed poorly on the Background class with an accuracy of only 0.0243, essentially failing to recognize the latter. ProtoNet showed a similar pattern, performing well on the Rain and Smash classes, with respective accuracies of 0.8419 and 0.8864, but yielded zero accuracy on the other two. This indicates its performance is fragile and class-selective under deployment shifts.
Table 6: Per-Class Accuracy (%) and Mean on OSDG1.
Model Background Climb Rain Smash Mean XGBoost $0.3783$ $0.0000$ $0.0081$ $0.7674$ $0.1463$ SVM $0.3783$ $0.0000$ $0.0000$ $0.9069$ $0.1615$ KNN $0.0243$ $0.1621$ $0.9919$ $0.5116$ $0.4725$ 1D-CNN $0.0000$ $0.0000$ $0.5945$ $0.9999$ $0.4420$ 2D-CNN $0.9999$ $0.9999$ $0.9999$ $0.0406$ $0.5914$ LSTM $0.0000$ $0.6511$ $0.5405$ $0.0162$ $0.1524$ Transformer $0.0079$ $0.2558$ $0.6486$ $0.2276$ $0.1951$ ProtoNet $0.0000$ $0.0000$ $0.8419$ $0.8864$ $0.4320$ MAML $0.4458$ $0.9354$ $0.5054$ $0.0400$ $0.4816$ ANIL $0.6258$ $0.6333$ $0.3100$ $0.1358$ $0.4262$ R2D2 $0.5681$ $0.9375$ $0.3333$ $0.0641$ $0.4804$ FEAT $0.9341$ $0.2730$ $0.1758$ $0.5800$ $0.4907$ DUPLE (Ours) $0.4208$ $0.3983$ $0.9891$ $0.7408$ $0.6372$
In the Climb test, MAML and R2D2 achieved high accuracy of 0.9354 and 0.9375 respectively, but their accuracy dropped sharply in the Smash test. The DUPLE model avoided catastrophic failures on any single class, achieving the best average accuracy of 0.6372. It maintained a high accuracy of 0.9891 in the Rain test, performed stably in the Smash test with an accuracy of 0.7408, and kept its accuracy within acceptable ranges in both the Background and Climb tests. This pattern suggests that the DUPLE model sacrifices some peak performance on simpler classes in exchange for significant improvements on more complex classes, which is crucial for handling data with domain drift issues.
Table 7: Per-Class Accuracy (%) and Mean on OSDG2.
Model Background Climb Impact Mean
XGBoost $0.3649$ $0.9999$ $0.6899$ $0.6440$
SVM $0.2499$ $0.9999$ $0.8099$ $0.6459$
KNN $0.2649$ $0.9007$ $0.8399$ $0.6384$
1D-CNN $0.7549$ $0.2595$ $0.8599$ $0.6723$
2D-CNN $0.9149$ $0.9847$ $0.1749$ $0.6534$
LSTM $0.9549$ $0.0305$ $0.9799$ $0.7363$
Transformer $0.9199$ $0.6183$ $0.3899$ $0.6459$
ProtoNet $0.6770$ $0.8959$ $0.6775$ $0.7501$
MAML $0.3120$ $0.9429$ $0.8483$ $0.7011$
ANIL $0.8159$ $0.6301$ $0.5043$ $0.6501$
R2D2 $0.8440$ $0.8514$ $0.5247$ $0.7400$
FEAT $0.9641$ $0.8019$ $0.2698$ $0.6786$
DUPLE (Ours) $0.7292$ $0.9861$ $0.7432$ $0.8195$
On OSDG2. Table 7 shows that the domain gap is narrowing, but similar trends still exist. Some deep models perform well in one class but poorly in another; for example, 2D-CNN performs well in background and Climb tests but poorly in Impact tests; FEAT performs well in background tests but poorly in Impact tests; LSTM is close to the upper limit in Impact tests but has almost zero accuracy in Climb tests. Meta-learners generally perform more evenly; ProtoNet and R2D2 provide reasonable coverage across all classes. DUPLE achieves the highest macro-average accuracy of 0.8195, with Climb accuracy close to the upper limit at 0.9861, and strong performance in background and Impact tests at 0.7292 and 0.7432, respectively. This indicates excellent cross-class robustness rather than over-optimization of a single class.
@@
In summary, on both datasets, the primary failure mode of the baseline model was single-class collapse caused by domain shift; accuracy was very high for one or two classes, while accuracy for other classes dropped to near zero. DUPLE’s dual-domain statistically guided prototyping reduces this disparity, improves accuracy for the worst-case class, and results in a more balanced and robust overall model performance, particularly for domain-shifted data. This is a valuable implementation for cross-deployment identification.
@@ 6 Conclusion
We propose DUPLE, a statistically guided dual-domain meta-learning framework designed specifically for event recognition in cross-deployment Distributed Fiber Optic Sensing (DFOS). DUPLE combines multi-prototype learners in the time and frequency domains with a statistically guided network (SGN) and a collaborative decision module (CDM). This design directly addresses three pain points of DFOS: domain shift due to fiber deployment methods, scarcity of new site labels, and intra-class diversity. The model learns expressive class representations and adjusts decisions during the inference phase without requiring target domain labels.

### v2

@@ 5.4 Comparison Experiments by Category
While overall accuracy provides a general measure of performance, it can obscure critical failures in specific classes, particularly under the imbalance induced by domain shifts. Therefore, we conduct a fine-grained analysis to determine whether models maintain reliable detection capabilities across all intrusion types or merely overfit to majority classes.
Analysis of OSDG1. As detailed in Table 12, standard deep learning baselines on OSDG1 suffer from catastrophic mode collapse. Notably, the 1DCNN failed to identify a single sample from the Background or Climb classes, yielding 0.00% accuracy in both cases, likely due to overfitting on the Smash class where it reached 76.87%. Even robust Domain Generalization (DG) methods exhibited significant instability; while GroupDRO performed well on Rain and Smash, it struggled severely with the Background class, achieving an accuracy of only 18.60%. This indicates that optimizing for worst-case groups does not fully resolve feature ambiguity in this domain.
Meta-learning baselines also demonstrated severe class selectivity. ProtoNet performed adequately on Background samples but its detection rate for Smash events dropped to a negligible 4.62%. These disjointed decision boundaries are visually corroborated by the confusion matrices in Figure 7. Specifically, GroupDRO (Figure 7e) shows substantial confusion between Background and Rain, while ProtoNet (Figure 7f) fails to establish a clear diagonal for the Smash class.
Figure 7: Per-class performance analysis on OSDG1 and OSDG2. (a)-(d) Precision and F1-scores. DUPLE is compared against representative baselines: GroupDRO (DG), ProtoNet, and FEAT (Meta-learning). While GroupDRO outperforms standard baselines, it shows instability on challenging classes (e.g., Rain in OSDG1). DUPLE (red) maintains robust performance across all categories. (e)-(l) Normalized confusion matrices. On OSDG1, GroupDRO (e) shows significant misclassification between Background and Rain. On OSDG2, GroupDRO (i) improves separation compared to metric-based methods like ProtoNet (j) but still confuses Impact with Background. DUPLE (l) exhibits the sharpest diagonal structure, indicating effective adaptation to target domains.
In contrast, the proposed DUPLE framework successfully mitigated these extreme disparities. DUPLE maintained balanced detection capabilities across all categories. Even on the most challenging Rain class, it retained a viable accuracy of 52.13%, a marked improvement over the near-zero failures observed in baseline methods. As illustrated in Figure 7(h), DUPLE recovers the diagonal structure of the confusion matrix, securing a consistent performance range of 52% to 69% across all four classes. This demonstrates that DUPLE prioritizes global reliability over peak performance on easier categories.
Analysis of OSDG2. Although the domain gap narrows on OSDG2, baseline models continued to exhibit distinct failure modes. Coral achieved perfect detection on Impact samples but sacrificed performance on the Climb class, where accuracy fell to 34.68%. Similarly, GroupDRO was heavily skewed, identifying 98.72% of Background samples but correctly classifying only 58.96% of Climb instances.
Among meta-learners, DUPLE demonstrated superior decision refinement. Quantitatively, DUPLE outperformed both GroupDRO and ProtoNet on the difficult Climb class with a leading accuracy of 86.41%, while maintaining strong performance on Background at 71.60% and Impact at 74.35%. The visual comparison between Figure 7(i) and Figure 7(l) further highlights this advantage: while GroupDRO misclassifies a significant portion of Impact samples as Background, DUPLE reduces this error and exhibits sharper diagonal dominance, indicating more precise decision boundaries.
@@
In summary, the primary failure mode of baseline models across both datasets was single-class collapse caused by domain shift. DUPLE’s dual-domain statistically guided prototyping strategy effectively reduces this class-wise disparity, ensuring that performance remains viable even for the worst-case categories.
@@ 6 Conclusion
We present DUPLE, a prototype-based meta-learning framework for cross-deployment DFOS recognition under scarce target labels. The key idea is to construct deployment-tolerant class representations by exploiting complementary time–frequency evidence and adjusting predictions according to sample-specific statistics. Concretely, DUPLE (i) adopts multi-prototype representations to capture intra-class heterogeneity, (ii) infers domain reliability from raw statistical cues, and (iii) performs query-adaptive prototype aggregation at inference. Experiments on OSDG1 and OSDG2 show consistent improvements over strong baselines and, importantly, mitigate class-wise worst-case degradations under deployment shifts, leading to more dependable recognition across all activity categories. Ablation and analysis further verify that these design choices provide complementary gains rather than redundant complexity.